OpenStudy (anonymous):

help! how do i simplify x-5/4x-20

OpenStudy (anonymous):

Divide each side by '9'. x = 2.222222222

OpenStudy (anonymous):

no i just need it simplified not the answer

OpenStudy (anonymous):

i got 1/4

OpenStudy (anonymous):

can you tell me the work though?

OpenStudy (anonymous):

Reorder the terms: 5x(20 + 4x) = 0 (20 * 5x + 4x * 5x) = 0 (100x + 20x2) = 0 Solving 100x + 20x2 = 0

OpenStudy (anonymous):

i dont think that is correct hmmmmmm

OpenStudy (espex):

Factor the denominator, this will allow you to cancel terms to simplify.